Mr Tui inherits a sum of $10,000 in Australia. He wants to invest this sum into three investments: Cooperate Bills, Cooperate Bonds and Treasury Bonds, with annual rates of 12%, 8% and 4% respectively. He wants his investment in Cooperate Bonds to be $500 more than the investment in Treasury Bonds. Use the Gauss Jordon Elimination method to find how much Mr Tui should allocate to each of his investments, in order to get an annual income of exactly $1,000.
Answer by jorel1380(3719)

Let m be the amount invested in Cooperate Bills and n be the amount in Cooperate Bonds. Then the amount in Treasury Bonds is n-500. Then:
m+n+n-500=10000
m+2n=10500
and
.12m+.08n+.04(n-500)=1000
.12m+.12n=1020 So:
m+2n=10500
.12m+.12n=1020
Solve for m and n
☺☺☺☺